this is answering simplifed questions with some mixed numbers single numbers and fractions I am sorry for the confusion
4/7*5
15/28 * 7/8
-3 1/5* 7/8

Answer by tommyt3rd(5050)

4/7*5 = 20/7 = 2 6/7

15/28 * 7/8 = 15/4 * 1/8 = 15/32

-3 1/5* 7/8 = -16/5 * 7/8 = -2/5 * 7/1 = -14/5 = -2 4/5

:)